# A theory of state-to-state transitions based on the framework of classical reaction dynamics

### v1

@@ IV.2 Binding kinetics
Table 4: Binding rate constants, $k_{\mathrm{on}}$, obtained from the present method and brute-force MD simulations. For the present method, the values of $k_{\mathrm{on}}$ are computed under the conditions $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$ and $\Delta R=1$$\mathrm{\AA }$. The statistical uncertainty is provided at 95% confidence interval.
@@
In this subsection, we discuss the application of the present method to molecular binding in the CH4/CH4, Na+/Cl-, and crown ether/K+ systems through the returning probability theory (Sec. II.5). $P_{\mathrm{RET}}\left(t\right)$ and $k_{\mathrm{ins}}$ involved in Eq. (43) are computed from the present method under the corresponding absorbing and reflecting boundary conditions. The value of $K^{*}$ is calculated based on the population ratio of the reactive state relative to an arbitrarily chosen unbound region including the standard state correction,[15, 26] with the US trajectories. For all cases, the state definition is based on the combination $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$, with a reactive state of width $\Delta R$. The lower bound of the reactive state corresponds to the boundary separating the bound region from the other regions, as described in the last paragraph of Sec. III.3. In the previous studies on protein-ligand binding[27] and membrane permeation[38], the appropriate width of the reactive state was determined such that the kinetic constants of interest are insensitive to the variation in the width.
@@
Fig. 8(b) shows the comparison of $k_{\mathrm{on}}$ between the present method and brute-force MD simulations. The values of $k_{\mathrm{on}}$ are also listed in Table 4. The present method estimates that $k_{\mathrm{on}}$ increases in the order of $\mathrm{CH_{4}/CH_{4}}>\mathrm{Na^{+}/Cl^{-}}>\mathrm{crown~ether/K^{+}}$. This ordering is consistent with that predicted from the brute-force MD simulations. Moreover, for all systems, the deviation of $k_{\mathrm{on}}$ from the values obtained by brute-force MD simulations is within $\sim 40\%$, indicating that the present method is useful for semi-quantitative ranking of binding kinetics with reduced computational cost compared to the brute-force approach.
@@ V Conclusion
In this study, a new method was developed to describe the time evolution of the populations of distinct states in a system of interest by means of molecular dynamics (MD) simulations. Based on mathematical techniques developed in the field of diffusion-influenced reaction theory, a Liouville equation for the probability density of each state was derived, in which the influx from and efflux to neighboring states are explicitly incorporated. Subsequently, a Markov approximation for the processes of crossing the boundaries separating the states was introduced, resulting in a tractable set of integral equations that describe the population of each state. Since the time-dependent quantities involved in the derived equations for a given state are connected to only a few states near the state, these quantities can be evaluated through the short-timescale MD simulations. We also presented a scheme for introducing absorbing and reflecting boundaries, making the method compatible with the returning probability theory for evaluating binding rate constants.
@@
The present method was applied to the binding and unbinding of the CH4/CH4, Na+/Cl-, and 18-crown-6-ether (crown ether)/K+ in water. For the unbinding processes, the time constants, $\tau_{\mathrm{off}}$, obtained from the present method were satisfactorily close to those from the brute-force MD simulations, with deviations of less than approximately 10%. Remarkably, for the crown ether/K+ system, the present method reproduced an unbinding timescale exceeding $10^{2}$ ns from a number of short-timescale (1–2 ns) simulations. Regarding the binding kinetics, the ordering of binding rate constants, $k_{\mathrm{on}}$, predicted from the present method was the same as that from the brute-force MD simulations. Furthermore, the deviation of $k_{\mathrm{on}}$ from the values obtained by the brute-force MD simulations was within $\sim$40%, enabling semi-quantitative ranking of the binding kinetics using MD simulations with timescales much shorter than those required for the brute-force MD approach.
@@
As demonstrated by the applications in the present study, the present method allows reliable estimation of the rate constants for binding and unbinding kinetics from the short-timescale MD simulations. Thus, complex protein–ligand binding systems, such as trypsin/benzamidine, Janus kinase (JAK)/inhibitors, threonine–tyrosine kinase (TTK)/inhibitors, and heat shock protein 90 (Hsp90)/inhibitors, for which the binding and unbinding rate constants have been experimentally measured,[43] are potential targets of the present method for elucidating the underlying binding mechanisms with quantitative assessment. To further expand the applicability of the present method, on the other hand, reducing the number of required trajectories is important. In the Markovian milestoning with Voronoi tessellations (MMVT) method,[68] the efficiency of sampling transition events is improved by introducing reflecting boundaries between states into MD simulations. Incorporating such simulation techniques could help reduce the computational cost of the present method. We anticipate that the present method and its extensions will be useful for unveiling the mechanisms underlying a variety of molecular kinetic processes in complex systems.

### v2

@@ IV.2 Binding kinetics
Table 4: Binding rate constants, $k_{\mathrm{on}}$, obtained from the present method (IEPDYN) and brute-force MD simulations. For the IEPDYN method, the values of $k_{\mathrm{on}}$ are computed under the conditions $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$ and $\Delta R=1$$\mathrm{\AA }$. The statistical uncertainty is provided at 95% confidence interval.
@@
In this subsection, we discuss the application of the IEPDYN method to molecular binding in the CH4/CH4, Na+/Cl-, and crown ether/K+ systems through the returning probability theory (Sec. II.5). $P_{\mathrm{RET}}\left(t\right)$ and $k_{\mathrm{ins}}$ involved in Eq. (44) are computed from the IEPDYN method under the corresponding absorbing and reflecting boundary conditions. The value of $K^{*}$ is calculated based on the population ratio of the reactive state relative to an arbitrarily chosen unbound region including the standard state correction,[15, 26] with the US trajectories. For all cases, the state definition is based on the combination $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$, with a reactive state of width $\Delta R$. The lower bound of the reactive state corresponds to the boundary separating the bound region from the other regions, as described in the last paragraph of Sec. III.3. In the previous studies on protein-ligand binding[27] and membrane permeation[38], the appropriate width of the reactive state was determined such that the kinetic constants of interest are insensitive to the variation in the width.
@@
Fig. 8(b) shows the comparison of $k_{\mathrm{on}}$ between the IEPDYN method and brute-force MD simulations. The values of $k_{\mathrm{on}}$ are also listed in Table 4. The IEPDYN method estimates that $k_{\mathrm{on}}$ increases in the order of $\mathrm{CH_{4}/CH_{4}}>\mathrm{Na^{+}/Cl^{-}}>\mathrm{crown~ether/K^{+}}$. This ordering is consistent with that predicted from the brute-force MD simulations. Moreover, for all systems, the deviation of $k_{\mathrm{on}}$ from the values obtained by brute-force MD simulations is within $\sim 40\%$, indicating that the IEPDYN method is useful for semi-quantitative ranking of binding kinetics with reduced computational cost compared to the brute-force approach.
@@ V Conclusion
In this study, we developed a new method, termed the integral-equation formalism of population dynamics (IEPDYN), to describe the time evolution of populations of distinct states in a system of interest using molecular dynamics (MD) simulations. Based on mathematical techniques developed in the field of diffusion-influenced reaction theory, a Liouville equation for the probability density of each state was derived, in which the influx from and efflux to neighboring states are explicitly incorporated. Subsequently, a Markov approximation for the processes of crossing the boundaries separating the states was introduced, resulting in a tractable set of integral equations that describe the population of each state. Since the time-dependent quantities involved in the derived equations for a given state are connected to only a few states nearby, these quantities can be evaluated through the short-timescale MD simulations. We also presented a scheme for introducing absorbing and reflecting boundaries, making the method compatible with the returning probability theory for evaluating binding rate constants.
@@
The IEPDYN method was applied to the binding and unbinding of the CH4/CH4, Na+/Cl-, and 18-crown-6-ether (crown ether)/K+ in water. For the unbinding processes, the time constants, $\tau_{\mathrm{off}}$, obtained from the IEPDYN method were satisfactorily close to those from the brute-force MD simulations, with deviations of less than approximately 10%. Remarkably, for the crown ether/K+ system, the IEPDYN method reproduced an unbinding timescale exceeding $10^{2}$ ns from a number of short-timescale (1–2 ns) simulations. Regarding the binding kinetics, the ordering of binding rate constants, $k_{\mathrm{on}}$, predicted from the IEPDYN method was the same as that from the brute-force MD simulations. Furthermore, the deviation of $k_{\mathrm{on}}$ from the values obtained by the brute-force MD simulations was within $\sim$40%, enabling semi-quantitative ranking of the binding kinetics using MD simulations with timescales much shorter than those required for the brute-force MD approach.
@@
The types of kinetic quantities that can be computed using the IEPDYN method, such as the mean first passage time (MFPT), are the same as those from the Markov state model (MSM). Both methods eanble us to compute these quantities using an identical set of state definitions. In contrast to the MSM approach, however, the IEPDYN method does not require the introduction of a coarse-grained timescale, commonly referred to as the lag time. This feature enables the computation of kinetic quantities without suffering from lag-time dependence. As demonstrated by the applications in the present study, the IEPDYN method allows reliable estimation of the rate constants for binding and unbinding kinetics from the short-timescale MD simulations. Thus, complex protein–ligand binding systems, such as trypsin/benzamidine, Janus kinase (JAK)/inhibitors, threonine–tyrosine kinase (TTK)/inhibitors, and heat shock protein 90 (Hsp90)/inhibitors, for which the binding and unbinding rate constants have been experimentally measured,[43] are potential targets of the IEPDYN method for elucidating the underlying binding mechanisms with quantitative assessment. To further expand the applicability of the IEPDYN method, on the other hand, reducing the number of required trajectories is important. In the Markovian milestoning with Voronoi tessellations (MMVT) method,[68] the efficiency of sampling transition events is improved by introducing reflecting boundaries between states into MD simulations. Incorporating such simulation techniques could help reduce the computational cost of the IEPDYN method. We anticipate that the IEPDYN method and its extensions will be useful for unveiling the mechanisms underlying a variety of molecular kinetic processes in complex systems.